OpenStudy (anonymous):

Multiply (6+√5)(4-2√5) Answers: 24-10√5 14-16√5 24-2√5 14-8√5

OpenStudy (angela210793):

24-12sqr5+4sqr5-10=14-8sqr5=14-8√5